# Shift orbifolds, decompactification limits, and lattices

### v1

@@ 4.3 Decompactification limits and T-dualities
Starting from a given point $(r,{{A}})$ in $\mathfrak{M}_{1,17}$, we let the radius $r$ decrease until we reach the boundary $r^{2}+\tfrac{1}{2}{{A}}\cdot{{A}}=1$ of the Narain moduli space. To stay inside $\mathfrak{M}_{1,17}$, we must perform a T-duality—a factorized duality, followed by a lattice shift by some $\boldsymbol{q}$ and a Weyl symmetry $R$ that brings the new Wilson line parameter back inside the fundamental alcove of the gauge moduli space. In this T-dual frame, the new moduli
@@
Our choice of $R$ and $\boldsymbol{q}$ is such that (4.21) is inside the fundamental alcove. At this point, there are three possibilities, depending on the value of ${{A}}$.
@@
    If ${{A}}$ is such that $\tfrac{2}{{{A}}\cdot{{A}}}{{A}}$ is a lattice vector, then $\boldsymbol{q}=-\tfrac{2}{{{A}}\cdot{{A}}}{{A}}$ and (4.21) vanishes. We perform an additional factorized duality. The new Wilson line parameter ${{A}}^{\prime\prime}$ goes to zero when $r\to 0$, while $r^{\prime\prime}$ runs off to infinity. Hence, we obtain a dual picture of the initial theory in which $r\to 0$ appears as a decompactification limit.
@@
    When ${{A}}$ is equal to ${\mathfrak{m}}$, then $\boldsymbol{q}=0$ and (4.21) equals ${\mathfrak{m}}$. We can use the Ginsparg map (4.2) to get to a dual frame, with moduli $r^{\prime\prime}$ and ${{A}}^{\prime\prime}$ obtained from (4.19). When $r\to 0$, the radius $r^{\prime\prime}$ goes to infinity: this is a decompactification limit, albeit of the other heterotic string.
@@
Applying the above procedure to the Wilson line parameter (4.4) will identify the $r\to 0$ decompactification limit with one of the two ten-dimensional heterotic strings. In fact, we can forget about the nine-dimensional description and rephrase this method at the level of the lattice $\Lambda$. We start with a given shift vector $s={\boldsymbol{\sigma}}/k$ associated with a cyclic shift orbifold of order $k$. Since translations of $s$ by lattice vectors do not affect the phase by which the orbifold acts on vertex operators, we can take $s$ to be in the maximal torus of the gauge group. Moreover, instead of $s$, we could pick any Weyl rotated shift vector $R(s)$—this amounts to working in a T-dual frame. Hence, without loss of generality, we can restrict to a shift vector inside the fundamental alcove:
@@
achieving the proof of the statement (4.25). The convergence of the sequence $\{s_{\text{i}}\}$ follows: if at the i-th step the vector $s_{\text{i}}={\boldsymbol{\sigma}}_{\text{i}}/k_{\text{i}}$ is not equal to 00 or ${\mathfrak{m}}$, then at the next step we have $s_{\text{i}+1}={\boldsymbol{\sigma}}_{\text{i}+1}/k_{\text{i}+1}$ with $k_{\text{i}+1}$ a positive integer strictly smaller than $k_{\text{i}}$. If the sequence did not converge, we would obtain an infinite sequence of positive integers with $k>k_{1}>\dots>k_{\text{i}}>k_{\text{i}+1}>\dots$, which is of course impossible. Hence, we see that after a finite number of steps (smaller than $k$), $s_{\text{i}}$ equals 00 or ${\mathfrak{m}}$. The Weyl symmetries $R_{\text{i}}$ and lattice elements $\boldsymbol{q}_{\text{i}}$ allow us to reconstruct, in nine dimensions, a sequence of dualities leading to a frame in which $r\to 0$ appears as a decompactification limit.
@@
Using the above construction we can characterize for which values of ${{A}}$ the $r\to 0$ limit is a decompactification limit. This will happen if and only if the Wilson line parameter ${{A}}$ is rational—i.e. if it has rational coefficients when expanded in a basis of simple roots (or equivalently, of coweights).
When the sequence of T-dualities introduced at the beginning of this subsection converges, ${{A}}$ is related to 00 or ${\mathfrak{m}}$ by a finite number of lattice vectors shifts and Weyl transformations. As a consequence, such an ${{A}}$ must be rational. Conversely, consider an arbitrary rational Wilson line
@@
where $m_{i}$ and $n_{i}$ are coprime integers for all $i=1,\dots,16$. We define $k=\text{lcm}(n_{1},\dots,n_{16})$ as the least common multiple of the integers in the denominators. In terms of $k$, the Wilson line parameter takes the form ${{A}}={\boldsymbol{\sigma}}/k$ for some lattice vector ${\boldsymbol{\sigma}}$. Let us apply the algorithm introduced above to ${{A}}$. We obtain a sequence of Wilson line parameters $\{{{A}}_{\text{i}}\}_{\text{i}\in\mathbb{N}}$, related to the initial one by a chain of T-dualities parametrized by $\boldsymbol{q}_{\text{i}}$ and $R_{\text{i}}$.
In order to prove the convergence of such a sequence, we relied on the condition ${\boldsymbol{\sigma}}\cdot{\boldsymbol{\sigma}}\in 2k{\mathbb{Z}}$—it allowed us to define $k_{1}$ at the first step by the equation ${\boldsymbol{\sigma}}\cdot{\boldsymbol{\sigma}}=2kk_{1}$. For a generic choice of rational Wilson line (4.30), this condition is not satisfied. Nevertheless, this is easily remedied by defining $k_{1}$ to be the numerator of $\frac{{\boldsymbol{\sigma}}\cdot{\boldsymbol{\sigma}}}{2k}\in\mathbb{Q}$ when expressed as an irreducible fraction. Doing so, $s_{1}$ can be recast in the form (4.25), and the corresponding lattice element ${\boldsymbol{\sigma}}_{1}$ satisfies the condition ${\boldsymbol{\sigma}}_{1}\cdot{\boldsymbol{\sigma}}_{1}\in 2k_{1}{\mathbb{Z}}$. The rest of the proof carries over, and the sequence of Wilson line converges to either 00 or ${\mathfrak{m}}$. This completes the analysis of $r\to 0$ limits of [4]: we see that in the presence of a rational Wilson line, the $r\to 0$ limit of the nine-dimensional heterotic theory is always a decompactification limit.
@@ 5.1 Niemeier lattices
[…]
@@
Starting from any of the twenty-three Niemeier lattices with roots $\Lambda_{\operatorname{\mathfrak{g}}}$, the Weyl vector can be used to construct the Leech lattice as a cyclic shift orbifold of $\Lambda_{\operatorname{\mathfrak{g}}}$ of order $h$—the Coxeter number of $\Lambda_{\operatorname{\mathfrak{g}}}$. The twenty-three resulting descriptions of the Leech lattice are known as its “holy constructions” [25, 31], and each can be understood as a ${\mathbb{Z}}_{h}$ shift orbifold with shift vector $s=\frac{{\boldsymbol{\rho}}}{h}$, which, as a result of (5.3), satisfies the orbifold consistency condition $s\cdot s\in\frac{2}{h}{\mathbb{Z}}$. Setting $\zeta=-1/h$, the invariant lattice takes the form
@@ 5.2 Orbifolds of the Leech CFT
[…]
@@
$\boldsymbol{f}^{\prime}$ $=k^{\prime}\left({{\boldsymbol{e}}}-\tfrac{s^{\prime}\cdot s^{\prime}}{2}{{\boldsymbol{e}}}^{\ast}-s^{\prime}\right)~{},$ with $\text{with }\qquad s^{\prime}$ $=\frac{k}{k^{\prime}}s+{\boldsymbol{q}}~{}.$ (5.25)

### v2

@@ 4.3 Decompactification limits and T-dualities
Starting from a given point (r,A)(r,{{A}}) in $\mathfrak{M}_{1,17}$, we let the radius $r$ decrease until we reach the boundary $r^{2}+\tfrac{1}{2}{{A}}\cdot{{A}}=1$ of the Narain moduli space. To stay inside $\mathfrak{M}_{1,17}$, we must perform a T-duality—a factorized duality, followed by a lattice shift by some $\boldsymbol{q}$ and a Weyl symmetry $R$ that brings the new Wilson line parameter back inside the fundamental alcove of the gauge moduli space. In this T-dual frame, the new moduli
@@
Our choice of $R$ and $\boldsymbol{q}$ is such that (4.21) is inside the fundamental alcove. At this point, there are three possibilities, depending on the value of A{{A}}.
@@
    If A{{A}} is such that $\tfrac{2}{{{A}}\cdot{{A}}}{{A}}$ is a lattice vector, then $\boldsymbol{q}=-\tfrac{2}{{{A}}\cdot{{A}}}{{A}}$ and (4.21) vanishes. We perform an additional factorized duality. The new Wilson line parameter ${{A}}^{\prime\prime}$ goes to zero when $r\to 0$, while $r^{\prime\prime}$ runs off to infinity. Hence, we obtain a dual picture of the initial theory in which $r\to 0$ appears as a decompactification limit.
@@
    When A{{A}} is equal to $\mathfrak{m}$}, then $\boldsymbol{q}=0$ and (4.21) equals $\mathfrak{m}$}. We can use the Ginsparg map (4.2) to get to a dual frame, with moduli $r^{\prime\prime}$ and ${{A}}^{\prime\prime}$ obtained from (4.19). When $r\to 0$, the radius $r^{\prime\prime}$ goes to infinity: this is a decompactification limit, albeit of the other heterotic string.
@@
Applying the above procedure to the Wilson line parameter (4.4) will identify the $r\to 0$ decompactification limit with one of the two ten-dimensional heterotic strings. In fact, we can forget about the nine-dimensional description and rephrase this method at the level of the lattice $\Lambda$. We start with a given shift vector $s={\boldsymbol{\sigma}}/k$ associated with a cyclic shift orbifold of order $k$. Since translations of $s$ by lattice vectors do not affect the phase by which the orbifold acts on vertex operators, we can take $s$ to be in the maximal torus of the gauge group. Moreover, instead of $s$, we could pick any Weyl rotated shift vector R(s)R(s)—this amounts to working in a T-dual frame. Hence, without loss of generality, we can restrict to a shift vector inside the fundamental alcove:
@@
achieving the proof of the statement (4.25). The convergence of the sequence $\{s_{\text{i}}\}$ follows: if at the i-th step the vector $s_{\text{i}}={\boldsymbol{\sigma}}_{\text{i}}/k_{\text{i}}$ is not equal to 0 or $\mathfrak{m}$}, then at the next step we have $s_{\text{i}+1}={\boldsymbol{\sigma}}_{\text{i}+1}/k_{\text{i}+1}$ with $k_{\text{i}+1}$ a positive integer strictly smaller than $k_{\text{i}}$. If the sequence did not converge, we would obtain an infinite sequence of positive integers with $k>k_{1}>\dots>k_{\text{i}}>k_{\text{i}+1}>\dots$, which is of course impossible. Hence, we see that after a finite number of steps (smaller than $k$), $s_{\text{i}}$ equals 0 or $\mathfrak{m}$}. The Weyl symmetries $R_{\text{i}}$ and lattice elements $\boldsymbol{q}_{\text{i}}$ allow us to reconstruct, in nine dimensions, a sequence of dualities leading to a frame in which $r\to 0$ appears as a decompactification limit.
@@
Using the above construction we can characterize for which values of A{{A}} the $r\to 0$ limit is a decompactification limit. This will happen if and only if the Wilson line parameter A{{A}} is rational—i.e. if it has rational coefficients when expanded in a basis of simple roots (or equivalently, of coweights).
When the sequence of T-dualities introduced at the beginning of this subsection converges, A{{A}} is related to 0 or $\mathfrak{m}$} by a finite number of lattice vectors shifts and Weyl transformations. As a consequence, such an A{{A}} must be rational. Conversely, consider an arbitrary rational Wilson line
@@
where $m_{i}$ and $n_{i}$ are coprime integers for all $i=1,\dots,16$. We define $k=\text{lcm}(n_{1},\dots,n_{16})$ as the least common multiple of the integers in the denominators. In terms of $k$, the Wilson line parameter takes the form ${{A}}={\boldsymbol{\sigma}}/k$ for some lattice vector $\boldsymbol{\sigma}$}. Let us apply the algorithm introduced above to A{{A}}. We obtain a sequence of Wilson line parameters $\{{{A}}_{\text{i}}\}_{\text{i}\in\mathbb{N}}$, related to the initial one by a chain of T-dualities parametrized by $\boldsymbol{q}_{\text{i}}$ and $R_{\text{i}}$.
In order to prove the convergence of such a sequence, we relied on the condition ${\boldsymbol{\sigma}}\cdot{\boldsymbol{\sigma}}\in 2k{\mathbb{Z}}$—it allowed us to define $k_{1}$ at the first step by the equation ${\boldsymbol{\sigma}}\cdot{\boldsymbol{\sigma}}=2kk_{1}$. For a generic choice of rational Wilson line (4.30), this condition is not satisfied. Nevertheless, this is easily remedied by defining $k_{1}$ to be the numerator of $\frac{{\boldsymbol{\sigma}}\cdot{\boldsymbol{\sigma}}}{2k}\in\mathbb{Q}$ when expressed as an irreducible fraction. Doing so, $s_{1}$ can be recast in the form (4.25), and the corresponding lattice element ${\boldsymbol{\sigma}}_{1}$ satisfies the condition ${\boldsymbol{\sigma}}_{1}\cdot{\boldsymbol{\sigma}}_{1}\in 2k_{1}{\mathbb{Z}}$. The rest of the proof carries over, and the sequence of Wilson line converges to either 0 or $\mathfrak{m}$}. This completes the analysis of $r\to 0$ limits of [4]: we see that in the presence of a rational Wilson line, the $r\to 0$ limit of the nine-dimensional heterotic theory is always a decompactification limit.
Let us note that the moduli space $\mathfrak{M}_{1,17}$ of the Narain CFT is a locally symmetric space: its simply-connected cover $\operatorname{Gr}(1,17)$ is the quotient of the Lie group $\operatorname{O{}}(1,17)$ by its maximally compact subgroup $\operatorname{O{}}(1)\times\operatorname{O{}}(17)$. As explained in [22, 23], the boundary of such spaces can be characterized in terms of rational parabolic subgroups of the underlying group—here, $\operatorname{O{}}(1,17)$. The analysis of [22, 23] applies more generally to heterotic compactifications on $T^{d}$ and the associated Narain CFTs, whose moduli space is $\mathfrak{M}_{d,d+16}$. It would be interesting to explore the relation between shift orbifolds and decompactification limits in this lower dimensional setting, and to characterize in detail rational parabolic subgroups of $\operatorname{O{}}(d,d+16)$ and their T-duality orbits.
@@ 5.1 Niemeier lattices
where each factor denotes the root lattice associated to a simply laced Dynkin diagram. This sublattice has full rank $\operatorname{rk}\operatorname{\mathfrak{g}}_{1}+\dots+\operatorname{rk}\operatorname{\mathfrak{g}}_{M}=24$, and all of its components share the same Coxeter number $h$. The lattice $\Lambda$ can be reconstructed by gluing to $\mathfrak{g}$ appropriate elements of the discriminant group $\operatorname{\mathfrak{g}}^{\ast}/\operatorname{\mathfrak{g}}$. [^18] We will not need the detailed form of these glue vectors. Their expression can be found in the classic reference [32]. We denote the corresponding Niemeier lattice by $\Lambda_{\operatorname{\mathfrak{g}}}$. The gluing construction does not add any new root to the lattice, so the total number of roots of $\Lambda_{\operatorname{\mathfrak{g}}}$ is equal to 24h24\,h. The twenty-four Niemeier lattices are the nodes in figure 1 below.
@@
Starting from any of the twenty-three Niemeier lattices with roots $\Lambda_{\operatorname{\mathfrak{g}}}$, the Weyl vector can be used to construct the Leech lattice as a cyclic shift orbifold of $\Lambda_{\operatorname{\mathfrak{g}}}$ of order hh—the Coxeter number of $\Lambda_{\operatorname{\mathfrak{g}}}$. The twenty-three resulting descriptions of the Leech lattice are known as its “holy constructions” [27, 33], and each can be understood as a ${\mathbb{Z}}_{h}$ shift orbifold with shift vector $s=\frac{{\boldsymbol{\rho}}}{h}$, which, as a result of (5.3), satisfies the orbifold consistency condition $s\cdot s\in\frac{2}{h}{\mathbb{Z}}$. Setting $\zeta=-1/h$, the invariant lattice takes the form
@@ 5.2 Orbifolds of the Leech CFT
such that any ${\boldsymbol{p}}\in\Gamma$ is uniquely written as ${\boldsymbol{p}}=\text{w}{{\boldsymbol{e}}}+\text{n}{{\boldsymbol{e}}}^{\ast}+{{\boldsymbol{L}}}$, where $\text{w},\text{n}$ are integers and ${{\boldsymbol{L}}}\in\Lambda_{\text{L}}$. The Leech lattice can be obtained by considering the sublattice ${{\boldsymbol{e}}}^{\perp}\subset\Gamma$ of points $\boldsymbol{p}$} orthogonal to $\boldsymbol{e}$}} (those with $\text{n}=0$), and modding out this sublattice by $\boldsymbol{e}$}} (which allows to also set $\text{w}=0$). Hence, we trivially have the isomorphism
@@
$\boldsymbol{f}^{\prime}$ $=k^{\prime}\left({{\boldsymbol{e}}}-\tfrac{s^{\prime}\cdot s^{\prime}}{2}{{\boldsymbol{e}}}^{\ast}-s^{\prime}\right)~,$ $\text{with }\qquad s^{\prime}$ $=\frac{k}{k^{\prime}}s+{\boldsymbol{q}}~.$ (5.25)